OpenStudy (anonymous):

Find the coordinates of the centroid of the triangle formed by the x-axis, the y-axis and the line 12x+9y=36. Show work, and please help!

OpenStudy (anonymous):

SORRY i wish i could help...

OpenStudy (tkhunny):

Do you have a method in mind? Calculus? Geometry?

OpenStudy (anonymous):

Geometry.

OpenStudy (tkhunny):

Then you need medians. Two will do, but it's always fun to draw the third, just to see that it hit's the same spot.

OpenStudy (anonymous):

I understand that, but the question that I offered if what the question is. If I need the medians then, how would I find them per say through the line that is given?

OpenStudy (tkhunny):

12x + 9y = 36 ==> \(\dfrac{x}{3} + \dfrac{y}{4} = 1\) x-intercept is (3,0) y-intercept is (0,4) The horizontal leg of your Right Triangle is length 3. It's midpoint is \(\left(\dfrac{3}{2},0\right)\) Where does that leave us?

OpenStudy (anonymous):

Finding the slope and then solving for y which would then lead to elimination to finding the coordinates of the centroid?

OpenStudy (tkhunny):

The centroid lies on the line containing (0,4) and (3/2,0), because this is one of the trangle's medians. Your task it to find another median and see where the two intersect.

OpenStudy (anonymous):

Oh, I get it now. I got it from here. Thanks for your help. I appreciate it.